# Randomised benchmarking for universal qudit gates

### v1

@@ 4 Results
In this section, we describe the algorithm used to estimate the AGF of a rHDG gate set. By using a rHDG gate set and the randomised benchmarking algorithm, we extend randomised benchmarking to characterise universal gates. In practice, our scheme is useful to characterise single-qudit T gates and also controlled T gates; our scheme also characterises any diagonal unitary gate with order $\geq d$.
@@ 4.1 Randomised benchmarking and the real hyperdihedral group
The expression of the twirl $\mathcal{T}_{\mathcal{E}}$ and the quantities $\eta_{i}$ are consequences of Schur’s lemma [85] and our decomposition in Eq. (3.10). The projectors are explicitly constructed in Corollary 4.4. ∎
@@
with $A_{\varpi}\coloneqq\langle\!\langle\varpi|\mathcal{E}\Pi_{\mathbb{I}}|\varpi\rangle\!\rangle$$B_{\varpi}\coloneqq\langle\!\langle\varpi|\mathcal{E}\Pi_{\varpi}|\varpi\rangle\!\rangle$, and $\eta_{\varpi}$ given in Eq. (4.2). Note that the parameters $\eta_{\varpi}$ are ‘parts’ of the trace of the Pauli-Liouville representation of the noise. Therefore, by estimating these parameters we can estimate the trace of that Pauli-Liouville representation of the noise; this is a standard technique in randomised benchmarking schemes [64].
@@ 4.2 Experimental scheme
For the sake of concreteness, we consider the qutrit case. The primitive gates are
@@
the notation $(ij)$ refers to a permutation between the $i$-th and $j$-th elements in a list. We label the elements by a pair formed by a permutation $\sigma$ and three values $\bm{\alpha}=(\alpha_{0},\alpha_{1},\alpha_{2})$ with each $\alpha_{i}\in[9]$. The following matrices form the gate set:
@@
where each permutation $\sigma$ must be written in terms of $(12)$ and $(23)$ to only use the primitive gates.
We now illustrate a run with circuit depth $m=3$. The initial states required to prepare are
@@
$\begin{bmatrix}\omega_{9}^{3}&&\\ &\omega_{9}^{4}&\\ &&\omega_{9}^{8}\end{bmatrix}.$ (4.14)
Therefore, a run of the experiment with the previous sequence of gates corresponds to the following circuits:
@@
$C=\tilde{\gamma}((23),(7,8,8))\tilde{\gamma}((123),(0,5,7))\tilde{\gamma}((12),(8,1,4))\tilde{\gamma}(e,(3,4,8))$ (4.18)
are the gates in Eqs. (4.15) and (4.16). After the experimental procedure has been carried out, the resulting graph circuit depth vs averaged sequence fidelity is fitted to the curve in Eq. (4.7). We leave for §5 numerical results of our scheme and the statistical analysis of the number of shots and circuits required.
@@ 4.3 Characterisation of multi-qudit gates
In this subsection, we describe our scheme to characterise multiqudit gates. Whereas there are currently schemes to characterise Clifford gates in multilevel gates, there is a lack of methods for non-Clifford gates. Currently, there are at least three platforms with access to entangling gates is a superconducting platform [24, 4, 17, 94]; the fidelity achieved is $0.97$. Since the experimental scheme is the same as for the single-qudit case, we do not repeat the discussion.
@@
$\mathrm{CSUM}\ket{a,b}=\ket{a,a\oplus b};$
@@ 5 Numerics and construction of group elements for qutrits and ququarts
In this section I discuss the number of shots and circuits required for the estimate of the average gate fidelity for the qutrit and ququart cases. These results are to be used to determine the feasibility of our scheme. Our investigations are numerical because the variance is only possible to compute symbolically and not analytically [96].
@@
The structure of this section is as follows. First we discuss the numerical simulations and the interpretation of the outcomes. Then we present the results and comment on the conclusions that can be drawn from them. We conclude the section with a summary suggesting the experimental strategy suggested by the numerical results.
The simulations approximate the confidence intervals of the quantity estimated from a randomised benchmarking experiment using as gate set the rHDG. Our simulations use the central limit theorem for the shots; we approximated the distribution (corresponding to a randomly drawn circuit) with a normal distribution. For each circuit depth there is a normal distribution, corresponding to the sequence fidelity value resulting from the twirling procedure.
From each distribution we sample a sequence fidelity. This value of sequence fidelity is used to simulate the shots by using a Bernoulli random variable; the value of the sequence fidelity is the parameter of the Bernoulli’s distribution. In our numerical simulation we fix that number, which experimentally corresponds to the number of shots considered.
For a number of distinct random circuits, which is another experimental constraint, corresponding to a different value of sequence fidelity, we repeat the previous step, each time using a different value of the sequence fidelity. Averaging the resulting approximation of the sequence fidelity we obtain the final approximation of the sequence fidelity. The next part is the fitting procedure.
@@
Concluding the previous steps, we obtain a list of pairs. Each pair is formed by a circuit depth and the approximated sequence fidelity. For this list of values we draw a graph from which we fit using our expressions. For convenience, we only study the case for initial state |$\ket{0}$.
@@
Fitting the corresponding expression (Eq. (4.7)) to the graph explained in the previous paragraph then we estimate the parameter $\eta$; we denote the approximation $\tilde{\eta}$. We repeat the simulation step $10^{3}$ times, then compute $|\eta-\tilde{\eta}|$. From the values we compute quantiles to determine a pair error and confidence. The quantiles we use are 0.95, 0.99, and 1. Then, by reporting the number of shots and the number of randomly drawn circuits along the error and probability, we can provide estimates of the experimental statistical elements (shots and circuits) necessary for a sensible characterisation.
@@
We limited our search to 10 and 100 cycles to be a representative figure of an actual experiment; these numbers are also within the range of numbers has been reported in the literature [36]. In order to assess the spectrum of quality of gates, our simulation is parametrised by the average gate fidelity of the noise channel. We use a random channel [97] which is non-unital but CPTP. The results are presented as follows.
@@
Table 1:
@@
Table 2:
@@
Table 3:
@@
Table 4:
Each table should be read as follows. If the noise is known or it is suspected that has average gate fidelity in one of the values proposed, then, according to the confidence expected (we use 0.95, 0.999, and 1), the shots and randomly drawn circuits are displayed. For concreteness, we explain the reading of Tab. 1: fixing the number of shots and circuits as 100 and for a noise with fidelity $0.89$, the values resulting from the fitting of Eq. (4.7) have an error of $0.003$ with a frequency of $0.95$, error $0.006$ with frequency $0.999$, and error $0.008$ in any case.
For the ququart case, we get similar results:
@@
Table 5:
@@
Table 6:
@@
Table 7:
@@
In this paragraph, I discuss the best circuit depth strategy to obtain smaller errors in the estimate parameters $\eta_{\varpi}$. First, fix a channel with an average gate fidelity of $0.95$, which is a sensible value in current implementations. Then, estimate the parameter $\eta_{0}$ according to certain chosen circuit depths. The strategies chosen are: 1. $m\in[5]$, 2. close values $m\in[8]$, 3. spaced values $m\in\{5,10,15,20,\ldots,100\}$, and 4. $m\in\{2,4,8,16,32,64\}$. The best strategy is taking multiples of $5$ (strategy iii). The order of strategies is iii, ii, i, iv. Changing the spacing in strategy iii did not affect the result significantly.
@@
$T_{0}\coloneqq\operatorname{diag}(\omega_{9}^{2},\omega_{9}^{4},\omega_{9}^{3})\text{ and }T_{1}\coloneqq\operatorname{diag}(\omega_{9}^{2},\omega_{9}^{2},\omega_{9}^{5}).$ (5.1)
@@
$\langle\sigma_{x}\times\mathbb{I},\sigma_{z}\times\mathbb{I},\mathbb{I}\times\sigma_{x},\mathbb{I}\times\sigma_{z}\rangle,$ (5.3)
where $\sigma_{x}$ and $\sigma_{z}$ and the $X$ and $Z$ matrices for qubits. The T gate is $T\times\mathbb{I}$ with $T=\operatorname{diag}(1,\omega_{8})$.
@@ Appendix A Construction of group elements
$D_{1}(\operatorname{\mathfrak{o}}(d))\coloneqq\operatorname{diag}(\omega_{\operatorname{\mathfrak{o}}(d)},1,\ldots,1),$ (A.1)
@@ Appendix B Identifying the product of cyclic groups
${}_{d}\mathrm{rHDG}$ is a semidirect product of a symmetric group and a product of cyclic groups. This product of cyclic group depends on the matrix $T$. In this appendix, we clarify the process to identify the product of cyclic groups. This appendix is relevant for §3.
@@ C.2 Number of irreps in the Pauli-Liouville representation
note that $\chi_{\Gamma}(\sigma,\bm{\alpha})$ is a real number, thus I omit the modulus appearing in Theorem C.1. The next step is averaging ${\chi_{\Gamma}}^{2}$ (the square of $\chi_{\Gamma}(\sigma,\bm{\alpha})$) over every ${}_{d}\mathrm{rHDG}$ element and verifying that $\operatorname*{\mathbb{E}}_{\sigma,\bm{\alpha}}{\chi_{\Gamma}(\sigma,\bm{\alpha})}^{2}=3$. Then, using Theorem C.1, show that there are only three inequivalent irreps in the ${}_{d}\mathrm{rHDG}$.
First, I compute $\chi^{2}_{\Gamma}(\sigma,\alpha)$:
@@
To simplify Eq. (C.8e) I need to isolate the phases $\alpha_{i}$ appearing in the exponent of $\omega$. Only then I can compute the average over the phases. I decompose the cases $i\neq j$ and $u\neq v$ in Table 8.
Table 8: Table with the clasification of the phases.
@@
On Table 8 I describe the different combinations of phases that appear in the sum Eq. (C.8e). I decompose the sum-average in the second summand in the right-hand side of Eq. (C.8e):
@@
Note that each sum, except (C.8g), has in the exponent a sum of phases, with each phase different. Thus, averaging over the phases I obtain zero. Only the sum (C.8g) is non-zero:
@@
Note that the computations needed to arrived at the last equation (Eq. (C.8k)) are valid for any phase order greater than two. This shows my scheme is valid for any diagonal gate with order greater than two.
@@
Let $n_{i}\in\mathbb{Z}^{+}$ for $i\in[3]$. Then the only solution to the equation $\sum n_{i}^{2}=3$ is to have three non-zero $n_{i}$ and those non-zero are equal to $1$. Therefore, by Theorem C.1, there are three inequivalent irreps in the decomposition of the PL representation of the rHDG. ∎

### v2

@@ 4 Results
In this section, we describe the algorithm used to estimate the AGF of a rHDG gate set. By using a rHDG gate set and the standard randomised benchmarking [69], we extend randomised benchmarking to characterise universal gates. In practice, our scheme is useful to characterise single-qudit T gates and also controlled T gates; our scheme also characterises any diagonal unitary gate with order greater or equal than $d$.
@@ 4.1 Randomised benchmarking and the real hyperdihedral group
The expression of the twirl $\mathcal{T}_{\mathcal{E}}$ in Eq. (4.1) and the quantities $\eta_{i}$ are consequences of both Schur’s lemma [85] and our decomposition in Eq. (3.10). The projectors are explicitly constructed in Corollary 4.4. ∎
@@
with $\varpi\in\{0,+\}$,$A_{\varpi}\coloneqq\langle\!\langle\varpi|\mathcal{E}\Pi_{\mathbb{I}}|\varpi\rangle\!\rangle$$B_{\varpi}\coloneqq\langle\!\langle\varpi|\mathcal{E}\Pi_{\varpi}|\varpi\rangle\!\rangle$, and $\eta_{\varpi}$ given in Eq. (4.2). Note that the parameters $\eta_{\varpi}$ are ‘parts’ of the trace of the Pauli-Liouville representation of the noise. Therefore, by estimating these parameters we can estimate the trace of that Pauli-Liouville representation of the noise; this is a standard technique in randomised benchmarking schemes [64].
@@ 4.2 Experimental scheme
For definiteness, we consider the qutrit case. The primitive gates are
@@
the notation $(ij)$ refers to a permutation between the $i$-th and $j$-th elements in a list. We label the elements of the ${}_{d}\mathrm{rHDG}$ by a pair formed by a permutation $\sigma$ and three values $\bm{\alpha}=(\alpha_{0},\alpha_{1},\alpha_{2})$ with each $\alpha_{i}\in[9]$. The following matrices form the gate set:
@@
where each permutation $\sigma$ must be written as products of $(12)$ and $(23)$ so as to only use the primitive gates.
We now illustrate an example run, taking circuit depth $m=3$. The initial states required to prepare are
@@
$\begin{bmatrix}\omega_{9}^{3}&&\\ &\omega_{9}^{4}&\\ &&\omega_{9}^{8}\end{bmatrix},$ (4.14)
corresponding to the ${}_{d}\mathrm{rHDG}$ element $\gamma(e,(3,4,8))$. Therefore, a run of the experiment with the previous sequence of gates corresponds to the following circuits:
@@
$Q=\tilde{\gamma}((23),(7,8,8))\tilde{\gamma}((123),(0,5,7))\tilde{\gamma}((12),(8,1,4))\tilde{\gamma}(e,(3,4,8))$ (4.18)
are the gates in Eqs. (4.15) and (4.16). After the experimental procedure has been carried out, the resulting graph circuit depth vs averaged sequence fidelity is fitted to the curve of Eq. (4.7). We leave for §5 numerical results of our scheme and the statistical analysis of the number of shots and circuits required.
@@ 4.3 Characterisation of multi-qudit gates
In this subsection, we describe our scheme to characterise multiqudit gates. Whereas there are currently schemes to characterise Clifford gates in multilevel gates, there is a lack of methods for non-Clifford gates. Currently, there are at least three platforms with access to entangling gates [24, 4, 17, 94]. Since the experimental scheme is the same as for the single-qudit case, we do not repeat the discussion.
@@
$\mathrm{CSUM}\ket{a,b}=\ket{a,a\oplus b};$ (4.19)
@@ 5 Numerics and construction of group elements for qutrits and ququarts
In this section we discuss the number of shots and circuits required for the estimate of the average gate fidelity for the qutrit and ququart cases. These results are to be used to determine the feasibility of our scheme. Our investigations are numerical because the variance of the sequence fidelity for the ${}_{d}\mathrm{rHDG}$ is currently only possible to compute symbolically (using Mathematica) and not analytically [96].
@@
The structure of this section is as follows. First we discuss the numerical simulations and the interpretation of their outcomes. Then we discuss the conclusions drawn from the resulting confidence intervals. We conclude the section with a summary proposing an experimental strategy supported by our results. Our simulations approximate the confidence intervals of the parameter $\eta_{0}$ estimated from a randomised benchmarking experiment using as gate set the ${}_{d}\mathrm{rHDG}$.
In a randomised benchmarking scheme there are two important quantities in the design of the experiment: the number of shots and the nubmer of circuits used. In this section, we estimate the optimal values for these experimental parameters based on numerical simulations. First, we describe our simulation linking it with an hypothetical experiment.
Given a circuit depth, we must choose a sequence of gates, referred to as a circuit. The value of the fidelity for the initial state under the action of the circuit is called the sequence fidelity. Thus, for a given circuit depth $m$, we have a list of possible sequences of gates; if the gate set has order $|\mathbb{G}|$, the number of possible sequences is $|\mathbb{G}|^{m}$.
In an experiment, only a sequence of 00s and $1$s is recorded, depending on whether the initial state is recovered or not: click or no click. The distribution of clicks or no clicks follows a Bernoulli distribution. Thus, for a given circuit, the sequence fidelity is the mean of a list of Bernoulli distribution. By the central limit theorem, we can approximate the mean of a list of Bernoulli distributions by a normal distribution. In turn, the distribution of the mean sequence fidelity, $\Pr(m;\ket{\psi},|\psi\rangle\!\langle\psi|,\bm{\mathcal{E}})$ in Eq. (4.7), is the mean of a list of normal distributions. Therefore, $\Pr(m;\ket{\psi},|\psi\rangle\!\langle\psi|,\bm{\mathcal{E}})$ is also a normal distribution, with the mean given by Eq. (4.7).
Therefore, our numerical simulation can be summarised in the following steps
  1. 1.
    Fix circuit depth $m$, number of shots $s$, and number of random circuits $r$.
  2. 2.
    Obtain the sequence fidelity corresponding to the twirl to the power $m$; as argued above, this is the mean of a normal distribution that we sample in a future step.
  3. 3.
    Obtain the variance of the sequence fidelity [97].
  4. 4.
    Randomly sample the normal distribution to obtain an approximate sequence fidelity $\check{p}$.
  5. 5.
    Compute a Bernoulli random variable parametrised by $\check{p}$. Use this Bernoulli random variable to sample a binary sequence of length $s$. The mean is the approximate sequence fidelity of an experiment: call it $p^{\prime}$.
  6. 6.
    Repeat the previous steps (ii-v) to average $r$ different values of $p^{\prime}$ that approximates $\Pr(m;\ket{\psi},|\psi\rangle\!\langle\psi|,\bm{\mathcal{E}})$.
@@
Concluding the previous steps, we obtain a list of pairs. Each pair is formed by a circuit depth and the approximated sequence fidelity. For this list of values we construct a graph from which we fit using Eq. (4.7), obtaining a parameter $\check{\eta}_{0}$. For convenience, we only study the case for initial state |$\ket{0}$.
@@
Fitting (according to Eq. (4.7)) the graph constructed as explained in the previous paragraph yields an estimate of the parameter $\eta_{0}$; we denote the approximation $\tilde{\eta}$. We repeat the simulation step $10^{3}$ times, then compute $|\eta-\tilde{\eta}|$. From these values we compute quantiles to determine an error and confidence. The quantiles we use, for the difference $|\eta_{0}-\tilde{\eta}_{0}|$, are 0.95, 0.99, and 1. Then, by reporting the number of shots and the number of randomly drawn circuits, paired with the error and confidence, we can provide estimates of the experimental statistical elements (shots and circuits) necessary for a sensible characterisation.
@@
We limited our simulations to 10 and 100 cycles so as to be a representative figure for an actual experiment; these numbers are also within the range of numbers reported in the literature [36]. In order to assess the spectrum of quality of gates, our simulation is parametrised by the average gate fidelity of the noise channel. We use a random channel [98] which is non-unital but CPTP. The results are presented as follows.
@@
Table 1: Table reporting the confidence values for the estimate of the parameter $\eta_{0}$. The noise corresponds to a randomly sampled channel with average gate fidelity $0.89$.
@@
Table 2: Table reporting the confidence values for the estimate of the parameter $\eta_{0}$. The noise model is a composition of a totally depolarising and an amplitude damping channel, which is a comprehensive Markovian and completely positive noise model. The value of fidelity corresponds to the parameters $0.01$ for the depolarising and amplitude damping channel.
@@
Table 3: Table reporting the confidence values for the estimate of the parameter $\eta_{0}$. The noise model is a composition of a totally depolarising and an amplitude damping channel, which is a comprehensive Markovian and completely positive noise model. The value of fidelity corresponds to the parameters $0.03$ for the depolarising and amplitude damping channel.
@@
Table 4: Table reporting the confidence values for the estimate of the parameter $\eta_{0}$. The noise model is a composition of a totally depolarising and an amplitude damping channel, which is a comprehensive Markovian and completely positive noise model. The value of fidelity corresponds to the parameters $0.05$ for the depolarising and amplitude damping channel.
Each table should be read as follows. If the noise is known or it is suspected to have an average gate fidelity value in one of the values on Tables 1-4 then, according to the confidence expected to obtain (we use 0.95, 0.999, and 1), the number of shots and circuits required to obtain such confidence intervals are displayed. For concreteness, we explain the reading of Tab. 1: fixing the number of shots and circuits to 100 and for a noise with fidelity $0.89$, the values resulting from the fitting of Eq. (4.7) have an error of $0.003$ with a frequency of $0.95$, error $0.006$ with frequency $0.999$, and error $0.008$ in any case. This could also be written as $\mathrm{Prob}(|\eta_{0}-\check{\eta}_{0}|>0.003)=1-0.95$, $\mathrm{Prob}(|\eta_{0}-\check{\eta}_{0}|>0.006)=1-0.999$, and $\mathrm{Prob}(|\eta_{0}-\check{\eta}_{0}|>0.008)=0$, respectively.
For the ququart case, we get similar results, as presented in Tables 5-7:
@@
Table 5: Table reporting the confidence values for the estimate of the parameter. In contradistinction with the qutrit case, the noise model corresponded to a randomly sampled channel with an average gate fidelity $0.90$.
@@
Table 6: Table reporting the confidence values for the estimate of the parameter. In contradistinction with the qutrit case, the noise model corresponded to a randomly sampled channel with an average gate fidelity $0.99$.
@@
Table 7: Table reporting the confidence values for the estimate of the parameter. In contradistinction with the qutrit case, the noise model corresponded to a randomly sampled channel with an average gate fidelity $0.95$.
@@
Strategy Error given a probability
circuit depth 0.95 0.999 1
i 0.002 0.003 0.004
ii 0.002 0.003 0.004
iii 0.003 0.004 0.006
iv 0.001 0.001 0.002
Table 8: Comparison of the strategies. The result is based on the confidence intervals. For the noise we used a depolarising channel composed with a phase-damping; the depolarising and dephasing parameters are set to $0.1$.
In this paragraph, we discuss the best circuit depth strategy to obtain smaller errors in the estimate parameters $\eta_{\varpi}$. The strategy is drawn from the results presented in Table 8. First, fix the noise to be a composition of a depolarising and phase-damping channels, which amounts for a comprehensive kind of noise present in current implementations. Then, estimate the parameter $\eta_{0}$ according to certain chosen circuit depths. The strategies chosen are: 1. $m\in\{2,4,8,16,32,64\}$. 2. $m\in[8]$, 3. $m\in[5]$, 4. spaced values $m\in\{5,10,15,20,\ldots,100\}$. The best strategy, resulting from the simulation and reported on Table 8, is taking multiples of $5$ (strategy iv). The order of strategies is iv, i, ii, iii. Changing the spacing in strategy iv did not improve the result significantly.
@@
$T_{0}\coloneqq\operatorname{diag}[\omega_{9}^{2},\omega_{9}^{4},\omega_{9}^{3}]\text{ and }T_{1}\coloneqq\operatorname{diag}[\omega_{9}^{2},\omega_{9}^{2},\omega_{9}^{5}].$ (5.1)
@@
$\langle\sigma_{x}\otimes\mathbb{I},\sigma_{z}\otimes\mathbb{I},\mathbb{I}\otimes\sigma_{x},\mathbb{I}\otimes\sigma_{z}\rangle,$ (5.3)
where $\sigma_{x}$ and $\sigma_{z}$ and the $X$ and $Z$ matrices for qubits. The T gate is $T\times\mathbb{I}$ with $T=\operatorname{diag}[1,\omega_{8}]$.
@@ Appendix A Construction of group elements
$D_{1}(\operatorname{\mathfrak{o}}(d))\coloneqq\operatorname{diag}[\omega_{\operatorname{\mathfrak{o}}(d)},1,\ldots,1],$ (A.1)
@@ Appendix B Identifying the product of cyclic groups
${}_{d}\mathrm{rHDG}$ is a semidirect product of a symmetric group and a product of cyclic groups. This product of cyclic groups depends on the matrix $T$. In this appendix, we clarify the process to identify the product of cyclic groups. This appendix is relevant for §3.
@@ C.2 Number of irreps in the Pauli-Liouville representation
note that $\chi_{\Gamma}(\sigma,\bm{\alpha})$ is a real number, thus we omit the modulus appearing in Theorem C.1. The next step is averaging ${\chi_{\Gamma}}^{2}$ (the square of $\chi_{\Gamma}(\sigma,\bm{\alpha})$) over every ${}_{d}\mathrm{rHDG}$ element and verifying that $\operatorname*{\mathbb{E}}_{\sigma,\bm{\alpha}}{\chi_{\Gamma}(\sigma,\bm{\alpha})}^{2}=3$. Then, using Theorem C.1, show that there are only three inequivalent irreps in the ${}_{d}\mathrm{rHDG}$.
First, we compute $\chi^{2}_{\Gamma}(\sigma,\alpha)$:
@@
To simplify Eq. (C.8e) we need to isolate the phases $\alpha_{i}$ appearing in the exponent of $\omega$. Only then we can compute the average over the phases. We decompose the cases $i\neq j$ and $u\neq v$ in Table 9.
Table 9: Table with the clasification of the phases of Eq. (C.8e).
@@
On Table 9 we describe the different combinations of phases that appear in the sum Eq. (C.8e). We decompose the sum-average in the second summand in the right-hand side of Eq. (C.8e):
@@
Note that each sum, except (C.8g), has in the exponent a sum of phases, with each phase different. Thus, averaging over the phases we obtain zero. Only the sum (C.8g) is non-zero:
@@
Note that the computations needed to arrived at the last equation (Eq. (C.8k)) are valid for any phase order greater than two. This shows our scheme is valid for any diagonal gate with order greater than two. ∎
@@
Let $n_{i}\in\mathbb{Z}^{+}$ for $i\in[3]$. Then the only solution to the equation $\sum n_{i}^{2}=3$ is to have three non-zero $n_{i}$ and those non-zero are equal to $1$. Therefore, by Theorem C.1, there are three inequivalent irreps in the decomposition of the Pauli-Liouville representation of the rHDG. ∎